OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 4 blocks, 10 blocks, and 16 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? (5 points) Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied.

OpenStudy (anonymous):

A recursive formula: "Tells you how to 'grow' a sequence from a given first term" f(1) = 4 f(n) = f(n-1)+6 f(3) = 16 f(25) = 16 + (22*6) = 148 blocks. p.s. you posted in the writing section instead of the math section ;)

OpenStudy (anonymous):

lmaoooo ohhh i just realized that smh but thank you so much

OpenStudy (anonymous):

your welcome! :)